OpenStudy (anonymous):

Change 3x + 2y = 10 to slope-intercept form. You must show all work

OpenStudy (anonymous):

y = mx + c where m is the gradient (or slope) and c is the y intercept. Get it in to this form. Good luck.

OpenStudy (anonymous):

waiiit no how do i do that????

OpenStudy (anonymous):

add, subtract, multiply and divide both sides until it looks like y = mx + c. Give it a go and let me know how far you get.

OpenStudy (lalaly):

solve for y first subtract 3x from both sides\[2y=10-3x\]then divide by 2\[y=\frac{10-3x}{2}\]\[y=\frac{-3}{2}x+5\]

OpenStudy (anonymous):

thank you lala for showing me the steps how to do

OpenStudy (lalaly):

anytime ;)

OpenStudy (anonymous):

lalaly, she'll never learn if you do it all for her. it's better to guide her to the answer rather than giving it to her.

OpenStudy (lalaly):

i showed her the steps, she can use them in another question

OpenStudy (anonymous):

You showed her how to do a specific question. Following your method is not a general method and can only be applied to other questions if you already know the general method. I disagree that you helped at all, other than completing her homework.

OpenStudy (lalaly):

thats ur opinion...

OpenStudy (anonymous):

And you provided a strong rebuttal. Lols. Good day, sir.

OpenStudy (anonymous):

first of all this is recreational. I found a worksheet online and i needed help. she showed me how to do it so you can excuse yourself from this thread strobe. Thank you lala

OpenStudy (anonymous):

I tried to teach you how to solve all problems of this kind, rather than one specific example. I'm glad to see my help went unnoticed and unappreciated. You won't go far in life with that attitude, I'm sorry to say. Good day to you too.

OpenStudy (anonymous):

Alright when you have a timed test and need answer fast people give it to you because obviously they know what its like... Oh well if she gave her the answer, you can learn it later.